OpenStudy (anonymous):

if f(x) = 3x-1and g(x)=x-1, find (f o g)

OpenStudy (anonymous):

u mean find f(g(x))??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(g(x))= f(x-1) = (3x-1)-1 = 3x-2

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

your welcome good luck :)